Calculate the approximate pressure of a 2.0 mol sample of gas at 10.0 ºC and a volume of 15 L?





1
Expert's answer
2021-04-13T08:04:30-0400

10.0oC = 273 + 10 = 283 K

According to the ideal gas law,


"p=\frac{nRT}{V}=\frac{2.0mol\times8.314J/mol\cdot{K}\times283K}{15L}=310kPa"


Answer: 310 kPa